# Luminosity predictions for the first three ionisation stages of W, Pt and Au to probe potential sources of emission in kilonova

### v1

@@ Abstract
A large number of -matrix calculations of electron impact excitation for heavy elements () have been performed in recent years for applications in fusion and astrophysics research. With the expanding interest in heavy ions due to kilonova (KN) events such as AT2017gfo and GRB230307A, this new data can be utilised for the diagnosis and study of observed KN spectra. In this work recently computed electron-impact excitation effective collision strengths are used, for the first three ionisation stages of tungsten (W, ), platinum (Pt, ) and gold (Au, ), to construct basic collisional radiative models tailored for the late stage nebular phases of KN. Line luminosities are calculated at a range of electron temperatures and densities and the strengths of these lines for a representative ion mass are compared. For the case of W iii, these optically thin intensities are additionally used to constrain the mass of this ion in both AT2017gfo and GRB230307A. Comparing with theoretical predictions of nucleosynthesis yields from neutron star merger simulations, broad agreement with the inferred ion masses of W is found. Furthermore, we highlight the value of W measurements by showing that the abundance of other groups of elements and outflow properties are constrained by exploiting theoretically motivated correlations between the abundance of W and that of lanthanides or third r-process peak elements.
@@
Since the observation of the neutron star merger AT2017gfo kilonova (KN) event in 2017 (Smartt et al., 2017; Pian et al., 2017), several theoretical groups have been engaged in the possible identification of spectral features originating from r-process heavy elements. To date there has been confirmation of a P Cygni feature of Sr ii at approximately 1m (Watson et al., 2019) and another P Cygni feature of Y ii at 760nm (Sneppen & Watson, 2023). Collisional data using the -matrix method for the corresponding transitions of these two features have since been calculated by Mulholland et al. (2024a). In addition, a broad emission feature at 2.1m has been suggested by (Hotokezaka et al., 2023; Gillanders et al., 2024) originating from a forbidden transition between the fine-structure levels of the ground configuration of Te iii. This interpretation has been supported by NLTE modelling by Hotokezaka et al. (2023) and more recently by Mulholland et al. (2024b) using newly computed atomic data. A similar feature in GRB230307A has also been investigated by Levan et al. (2024); Mulholland et al. (2024b); Gillanders et al. (2023).
@@
There are, however, more recent atomic data now available in the literature for the first three ionisation stages of W. In Smyth et al. (2018) Dirac -matrix evaluations for the electron-impact excitation of W i are provided, Dunleavy et al. (2022) compute similar quality -matrix data for W ii and more recently McCann et al. (2024) for W iii. Using calibrated energy levels from NIST these three publications provide accurate Einstein A-values and effective collision strengths for a wide variety of incident electron temperatures. In this paper NLTE collisional radiative modelling is performed using these new atomic data sets for W i, ii and iii, to compute photon luminosity predictions to further investigate the W iii identification proposed by Hotokezaka et al. (2022) for the 4.5m forbidden line. The observability of the expected 6.05m W ii feature is also explored, which shows similar luminosities to the 4.5m line when ion masses similar to that of W iii are incorporated into the model. Additionally, the mass estimates of W iii in AT2017gfo from Hotokezaka et al. (2022) are revised using the newly calculated -matrix data and a first mass estimate for W iii in GRB230307A is provided. The mass estimates for W iii are used to compare with theoretically predicted nucleosynthesis yields, the implications of the W measurements for the production of other heavy elements such as lanthanides, actinides and third-peak r-process elements can then be investigated. The mass estimate of W is also important to understand its production by the r-process. So far, it has been observed in only a few metal-poor stars (Roederer et al., 2022).
@@
The outline of the paper is as follows. In Section 2 the general theory of collisional radiative modelling and the calculation of optically thin luminosity emission spectra is reviewed. In Section 3 synthetic spectra at late-time KN relevant conditions are presented. In addition, representative luminosities are presented for the ten strongest lines of the first three ion stages of W, Pt and Au. Mass estimates of W iii in both AT2017gfo and GRB230307Aare made with the parameter space of electron density and temperature explored. Furthermore in Section 4, the implications of the mass measurements based on comparisons with hydrodynamical neutron-star merger simulations and nucleosynthesis calculations are discussed. Finally in Section 5 the paper is concluded with a summary and outlook.
@@ 3.1 Luminosities
The electron density of the KN transient at late epochs is expected to be considerably lower than that of the early times. However, accurate values are difficult to estimate since they will depend on the radioactive heating, composition and ejecta density structure, which are hard to precisely constrain. Therefore, we will consider a range of electron densities in our analysis, cm-3, which are in line with values adopted in previous studies (Hotokezaka et al., 2023; Levan et al., 2024), for modelling AT2017gfo and GRB230307A. Since these densities are below their typical critical-densities, M1 transitions among the fine-structure split terms of the ground configuration are expected to give rise to strong emission lines in the infrared region of the observed spectrum. All three charge states of W of interest in this publication have fine-structure transitions among their ground state, for W i the ground level has terms 5d46s2 5D0,1,2,3,4, for W ii 5d46s 6D1/2,3/2,5/2,7/2,9/2 and for W iii 5d4 5D0,1,2,3,4. W iii is of particular interest as it has a strong fine-structure transition coincident with the Spitzer space telescope 4.5m wavelength range.
@@
In Figure 2 a model spectrum of W i-iii is plotted, which features line profiles broadened by a Gaussian kernel to produce a luminosity density spectrum. A full-width-half-maximum of 0.093 (Gillanders et al., 2023) was assumed for all lines. The kernels are normalised such that an integral over each line gives the total luminosity specified in Table 1. The spectrum is calculated at cm-3 for two different electron temperatures roughly representative of nebular KN events. To investigate blending between the ions spectra, each ion is treated separately. Evidently the strongest feature is that of the two W iii lines at 4.5 m. Notably, at these broadening parameters these two lines present themselves as essentially a single Gaussian feature. At a similar mass, it is found again that the W i and ii lines at around 6.05 m are of comparable strength. Additionally, there will be considerable blending between these two ion stages due to the overlap of the Gaussian profiles as seen in Figure 2. Nonetheless, it is feasible that future JWST observations could reach these large wavelengths in the IR, and in principle observe features here. This combined with the proposed W iii identification may provide conclusive evidence of the element W in future KN like events.
Recent -matrix calculations by McCann et al. (2022), Bromley et al. (2023) and Gillanders et al. (2021) have provided new radiative and collisional excitation atomic data for the first three ionisation stages of Pt and Au. To check for any additional potential sources of KN emission from these species the photon luminosity calculations are repeated for Pt i - iii and Au i - iii using this new atomic data. In Table 2 the calculated luminosities in units ph s-1 and erg s-1 for the strongest ten lines of each Au ion species are presented, and in Table 3 the corresponding values for the three Pt ions. Again the photon luminosities are computed with a reference mass of M⊙, an electron temperature = 0.15 eV and an electron density of = 106 cm-3. Clearly the luminosity strengths, for all six ions of both Au and Pt, are orders of magnitude lower than those predicted for W in Table 1, but the strongest transitions are similarly those among the low lying levels. For completeness in Figures 3 and 4 the corresponding broadened luminosity density profiles as a function of wavelength are plotted, to visually display these strong features. Given the lack of discernible features at these wavelengths at the late times of AT2017gfo and GRB230307A for Pt and Au, this places weak constraints on the masses of these elements as previously stated by (Gillanders et al., 2021). It is therefore the case that higher quality observed data is required to make more conclusive estimates of the presence of the elements in future KN events, in addition to accurate theoretical data for representative species.
@@ 3.2 Ion-mass quantification for W
The late-time-Spitzer (Villar et al., 2018; Kasliwal et al., 2019) observation of AT2017gfo and the observation of GRB230307A (Levan et al., 2024) have produced interest in a potential feature at around m. The ions W iii and Se iii, which both have thermally accessible fine structure lines at around this wavelength. A first analysis was performed by Hotokezaka et al. (2022) using distorted-wave collision data calculated using the hullac code (Bar-Shalom et al., 2001). Here, a revised analysis of the parameter space for W iii using the newly published -matrix collisional data of McCann et al. (2024) for AT2017gfo is presented, and additionally a new analysis for GRB230307A. This new atomic data is expected to be more accurate at low temperatures due to the inclusion of resonances when calculating the collision strengths.
@@
For GRB230307A, the feature has been integrated to have a specific luminosity of erg s-1 (Gillanders et al., 2023). Enforcing this luminosity, and a late time electron density of cm-3 (Levan et al., 2024), at an electron temperature of 3000 K (Levan et al., 2024), the -matrix data of McCann et al. (2024) requires a W iii mass of M⊙. Given a total ejecta mass of M⊙ (Gillanders et al., 2023), this mass of W iii represents a fraction of around %. Without a quantitative treatment of ionization fraction, an elemental mass is difficult to quantify. Assuming reasonably similar fractions for the near neutral stages of W, it is perhaps feasible that up to % could potentially be present in the ejecta.
Figure 5 shows the synthetic W spectrum compared with the JWST spectrum of GRB230307A (Levan et al., 2024). The centroid and width of the m feature was fitted by Gillanders et al. (2023), which is also shown in Figure 5. The spectrum was produced using an electron temperature of 0.26 eV and an electron density of cm-3. A Gaussian kernel was again assumed with a full-width-half-maximum of (Gillanders et al., 2023). With a W iii mass of M⊙, a good match to the JWST data can be found assuming a continuum at this point of erg s-1 nm -1.
@@
With increasing temperature at fixed density it is naturally found that less mass of W iii is required to produce the feature. The green dashed line on the top panel of Figure 6 shows the implications of a temperature of 0.86 eV (temperatures this high have been suggested by the models of Pognan et al., 2022). In this case, a relatively low mass of M⊙ of W iii is required to produce the feature. However, at such temperatures significant emission should begin to appear in the m range from W ii and iii and at around m from W i. While this in principle could constrain the mass of W i, the excess emission expected from W iii that is not present in the observation perhaps prevents this from being a valid identification. In summary, for W iii to be responsible for the excess flux at around m at the rough density regime of cm-3, it is required that the electron temperature also remain relatively low at around say 2500 K. Conversely, at low electron temperatures - say at around the K blackbody temperatures reported by Gillanders et al. (2023) the feature requires unreasonably high masses of around M⊙ of W iii.
@@ 4 Implications of the limits on the ejected mass of W
In this section, we compare theoretical merger models based on hydrodynamical simulations and nuclear network calculations with the estimated masses of W iii derived in the previous section, i.e. for GW170817 and for GRB230307A. Observational uncertainties of these estimates are discussed in the previous section, e.g. assumptions about the temperature and the possibility of line blending. The merger models only provide the total amount of W without distinguishing different ionization states and are themselves connected with uncertainties, some of which are briefly addressed below. Despite these uncertainties this comparison reveals a broad consistency between theoretical predictions and the observations and highlights the particular usefulness of W measurements to constrain the properties of the merger outflow and the abundance of other elements.
@@
In Table 4 we list the results from a selection of neutron star merger models, drawn from Just et al. (2023) and Sneppen et al. (2024), which strive for a complete and consistent description of all mass ejection channels. The set consists of four models taken from Just et al. (2023), two for equal-mass binaries ("sym") and two with a binary mass ratio of ("asy"). In order to capture the effects of turbulent angular-momentum transport during the post-merger evolution, these models adopt a viscosity model (Shakura & Sunyaev, 1973). For each of the aforementioned two cases we consider one model with a fairly strong (“n1-a6”) and one model with relatively weak (“n10-a3”) viscosity. The models of Just et al. (2023) only involve relatively long-lived neutron-star remnants that collapse 100 ms after merger or later (cf. in Table 4). We supplement these with two models (one equal mass, one asymmetric) from Sneppen et al. (2024) having short-lived remnants (i.e. collapsing at 10 ms after merger, labelled "short"; see references for details). The simulations consider neutron stars with masses comparable to those in GW170817. The total amount of ejecta in these simulations is also broadly consistent with the estimated total ejecta mass of AT2017gfo, namely about 0.03–0.05 (e.g. Smartt et al., 2017). In addition to the initial binary parameters and merger dynamics, nucleosynthesis yields are also affected by nuclear-physics uncertainties, which are particularly significant for heavy, neutron-rich nuclei. To explore this sensitivity, we add two further models that adopt the DZ31 nuclear mass model (Duflo & Zuker, 1995) compared to the models in Just et al. (2023); Sneppen et al. (2024), which adopted the HFB21 model (Goriely et al., 2010) as well as the consistently derived and rates in Mendoza-Temis et al. (2015). These two models are chosen to have all other simulation parameters matching previous cases such that the influence of changing the nuclear physics input alone can be ascertained.
@@
Elemental abundances in a KN are generally time dependent, because the radioactive decay half-lives are comparable to the relevant observable timescales of hours to months. The dominating decay mode, beta-decay, maintains constant the isotopic abundances but modifies the elemental abundances. To exemplify this, we show in Figure 7 the time evolution of the relative W mass fraction resulting in the simulation “sym-n1-a6” with the HFB21 nuclear mass model along with the contributions from individual isotopes. At around 1 day, 187W decays, and the total mass fraction decreases. The decays of the other two isotopes 188W and 185W (with half lives of around 70 and 75 days, respectively) lead to a further reduction (see supplemental material of Wu et al., 2019), despite a minor compensation from the decays of 184Ta and 183Ta. As a consequence, the mass fraction of W decreases by a factor of from the first hours to around two months after merger. Notice that this decrease should occur in any r-process model that reproduces the isotopic abundance pattern observed in the solar system around the third r-process peak. Given that the observations of AT2017gfo and GRB230307A discussed in Section 3.2 were taken at 43 days and 29 days, respectively, after the merger, we base the following quantitative comparison on the mass fractions taken at 1 month, which are listed in Table 4.
@@
The predicted W masses scatter in the different merger models between M⊙ and M⊙ and are thus broadly consistent with the mass of W iii derived through the luminosity of the 4.5 feature in the spectra of AT2017gfo and GRB230307A, providing tentative evidence that the inference presented in Section 3.2 yields reasonable estimates. The agreement is particularly good for AT2017gfo. Although most of the models give W masses that are a factor of 2–3 smaller than our estimated M⊙ (cf. Section 3.2), the full spread within the models encompasses this value. Also, several of the simulations yield mass fractions for W, (W), that are close to the estimated fraction of 0.3 % (Section 3.2). The short-lived merger models tend to yield somewhat higher values of (W) in slightly better agreement with the mass-fraction estimate from Section 3.2, however, the total ejecta masses in these models are slightly too small to be compatible with AT2017gfo. Considering GRB230307A, the inferred W iii mass of M⊙ is up to one order of magnitude in excess of what the simulations predict, which were, however, originally set up to reproduce GW170817-like systems. At this point we cannot judge if this discrepancy is a deficiency of the approach adopted in this study keeping in mind, in particular, as noted in Section 3.2, that the ion mass inferred from observations may be an overestimate if line blending is occurring, or only an upper limit if the feature is dominated by another process. Alternatively, this discrepancy may originate from different physical properties of the outflow in GRB230307A compared to AT2017gfo. As discussed below, a higher W mass points to more neutron-rich ejecta, which may result from a binary with different masses than GW170817. Since the gravitational-wave signal of GRB230307A was not detected, the binary masses in this event are unknown. High ejecta masses and very neutron-rich outflows may also result from a neutron-star black-hole merger (e.g. Kyutoku et al., 2021), which may also be a possible progenitor of GRB230307A. The numbers in Table 4 also show very clearly that the nuclear physics input for the network calculation has a significant impact on the W production, with DZ31 yielding systematically higher W masses (compare columns 2 and 8 or columns 7 and 9, respectively). The uncertainties of the nuclear model may thus represent another possible contribution explaining the difference between the theoretically predicted abundance and the W iii mass inferred from observations. Obviously, the current set of merger simulations is not exhaustive, and other physical ingredients (e.g. a different nuclear equation of state in the simulation) may yield different W masses. Moreover, due to their high level of complexity, the hydrodynamical simulations still adopt a number of approximations (e.g. in the neutrino transport and treatment of small-scale turbulence) and may not be fully resolved numerically.
@@
The robust pattern for the ratio of the W mass to the total ejecta mass with below 0.2, , suggests the possibility to use a measured W mass as a proxy for the amount of neutron-rich ejecta, the theoretical prediction of which, however, is still affected by nuclear-physics uncertainties. Based on the values from Table 4, neutron-rich material should roughly amount to about 40 times (14 times) the inferred W mass for HFB21 (DZ31), i.e. to M⊙ (M⊙) for AT2017gfo and to M⊙ (M⊙) for GRB230307A. For this neutron-rich material, the measured W mass constraints the production of non thermal electrons by beta decay. We find that for times later than a month 188W represents the dominating contribution.
@@
As an estimate of the uncertainties in the analysis and interpretation, which are typically a factor of a few (see Section 3.2), in Figure 9 we show a green band corresponding to a factor of two variation in both directions and a horizontal green band illustrating the corresponding uncertainties in (using the same conversion as above). Despite the remaining uncertainties our study suggests that a significant amount of lanthanides was produced in GW170817 and GRB230307A. We note that our estimated range of the lanthanide mass fraction (neglecting the small contributions from actinides) is in good agreement with the mass fraction of lanthanides that, according to the analysis of Ji et al. (2019), would be required by the observations of metal-poor stars for mergers to represent the main source of -process elements. Interestingly, the high lanthanide mass fractions that we find exceed most of the estimates made based on the light curve of AT2017gfo in the literature surveyed by Ji et al. (2019).
Similarly, our work predicts that a sizable amount of third-peak -process elements of about was co-produced (see lower panel of Figure 9).

### v2

@@ Abstract
A large number of -matrix calculations of electron impact excitation for heavy elements () have been performed in recent years for applications in fusion and astrophysics research. With the expanding interest in heavy ions due to kilonova (KN) events such as AT2017gfo and AT2023vfi, this new data can be utilised for the diagnosis and study of observed KN spectra. In this work recently computed electron-impact excitation effective collision strengths are used, for the first three ionisation stages of tungsten (W, ), platinum (Pt, ) and gold (Au, ), to construct basic collisional radiative models tailored for the late stage nebular phases of KN. Line luminosities are calculated at a range of electron temperatures and densities and the strengths of these lines for a representative ion mass are compared. For the case of W iii, these optically thin intensities are additionally used to constrain the mass of this ion in both AT2017gfo and AT2023vfi. Comparing with theoretical predictions of nucleosynthesis yields from neutron-star merger simulations, broad agreement with the inferred ion masses of W is found. Furthermore, we highlight the value of W measurements by showing that the abundance of other groups of elements and outflow properties are constrained by exploiting theoretically motivated correlations between the abundance of W and that of lanthanides or third r-process peak elements. Based on simple estimates, we also show that constraints on the distribution of tungsten in the ejecta may be accessible through the line shape, which may also yield information on the neutron-star merger remnant evolution.
@@
Since the observation of the neutron-star merger AT2017gfo kilonova (KN) event in 2017 (Smartt et al., 2017; Pian et al., 2017), several theoretical groups have been engaged in the possible identification of spectral features originating from r-process heavy elements. To date there has been confirmation of a P Cygni feature of Sr ii at approximately 1m (Watson et al., 2019) and another P Cygni feature of Y ii at 760nm (Sneppen & Watson, 2023). Collisional data using the -matrix method for the corresponding transitions of these two features have since been calculated by Mulholland et al. (2024a). In addition, a broad emission feature at 2.1m has been suggested by (Hotokezaka et al., 2023; Gillanders et al., 2024) originating from a forbidden transition between the fine-structure levels of the ground configuration of Te iii. This interpretation has been supported by NLTE modelling by Hotokezaka et al. (2023) and more recently by Mulholland et al. (2024b) using newly computed atomic data. A similar feature in AT2023vfi has also been investigated by Levan et al. (2024); Mulholland et al. (2024b); Gillanders & Smartt (2024).
@@
There are, however, more recent atomic data now available in the literature for the first three ionisation stages of W. In Smyth et al. (2018) Dirac -matrix evaluations for the electron-impact excitation of W i are provided, Dunleavy et al. (2022) compute similar quality -matrix data for W ii and more recently McCann et al. (2024) for W iii. Using calibrated energy levels from NIST these three publications provide accurate Einstein A-values and effective collision strengths for a wide variety of incident electron temperatures. In this paper NLTE collisional radiative modelling is performed using these new atomic data sets for W i, ii and iii, to compute photon luminosity predictions to further investigate the W iii identification proposed by Hotokezaka et al. (2022) for the 4.5m forbidden line. The observability of the expected 6.05m W ii feature is also explored, which shows similar luminosities to the 4.5m line when ion masses similar to that of W iii are incorporated into the model. Additionally, the mass estimates of W iii in AT2017gfo from Hotokezaka et al. (2022) are revised using the newly calculated -matrix data and a first mass estimate for W iii in AT2023vfi is provided. The mass estimates for W iii are used to compare with theoretically predicted nucleosynthesis yields, the implications of the W measurements for the production of other heavy elements such as lanthanides, actinides and third-peak r-process elements can then be investigated. The mass estimate of W is also important to understand its production by the r-process. So far, it has been observed in only a few metal-poor stars (Roederer et al., 2022).
@@
The outline of the paper is as follows. In Section 2 the general theory of collisional radiative modelling and the calculation of optically thin luminosity emission spectra is reviewed. In Section 3 synthetic spectra at late-time KN relevant conditions are presented. In addition, representative luminosities are presented for the ten strongest lines of the first three ion stages of W, Pt and Au. Mass estimates of W iii in both AT2017gfo and AT2023vfi are made with the parameter space of electron density and temperature explored. Furthermore in Section 4, the implications of the mass measurements based on comparisons with hydrodynamical neutron-star merger simulations and nucleosynthesis calculations are discussed. Finally in Section 6 the paper is concluded with a summary and outlook.
@@ 3.1 Luminosities
The electron density of the KN transient at late epochs is expected to be considerably lower than that of the early times. However, accurate values are difficult to estimate since they will depend on the radioactive heating, composition and ejecta density structure, which are hard to precisely constrain. Therefore, we will consider a range of electron densities in our analysis, cm-3, which are in line with values adopted in previous studies (Hotokezaka et al., 2023; Levan et al., 2024), for modelling AT2017gfo and AT2023vfi. Since these densities are below their typical critical-densities, M1 transitions among the fine-structure split terms of the ground configuration are expected to give rise to strong emission lines in the infrared region of the observed spectrum. All three charge states of W of interest in this publication have fine-structure transitions among their ground state, for W i the ground level has terms 5d46s2 5D0,1,2,3,4, for W ii 5d46s 6D1/2,3/2,5/2,7/2,9/2 and for W iii 5d4 5D0,1,2,3,4. W iii is of particular interest as it has a strong fine-structure transition coincident with the Spitzer space telescope 4.5m wavelength range.
@@
In Figure 2 a model spectrum of W i-iii is plotted, which features line profiles broadened by a Gaussian kernel to produce a luminosity density spectrum. A full-width-half-maximum of 0.11 (Gillanders & Smartt, 2024) was assumed for all lines. The kernels are normalised such that an integral over each line gives the total luminosity specified in Table 1. The spectrum is calculated at cm-3 for two different electron temperatures roughly representative of nebular KN events. To investigate blending between the ions spectra, each ion is treated separately. Evidently the strongest feature is that of the two W iii lines at 4.5 m. Notably, at these broadening parameters these two lines present themselves as essentially a single Gaussian feature. At a similar mass, it is found again that the W i and ii lines at around 6.05 m are of comparable strength. Additionally, there will be considerable blending between these two ion stages due to the overlap of the Gaussian profiles as seen in Figure 2. Nonetheless, it is feasible that future JWST observations could reach these large wavelengths in the IR, and in principle observe features here. This combined with the proposed W iii identification may provide conclusive evidence of the element W in future KN like events.
Recent -matrix calculations by McCann et al. (2022), Bromley et al. (2023) and Gillanders et al. (2021) have provided new radiative and collisional excitation atomic data for the first three ionisation stages of Pt and Au. To check for any additional potential sources of KN emission from these species the photon luminosity calculations are repeated for Pt i - iii and Au i - iii using this new atomic data. In Table 2 the calculated luminosities in units ph s-1 and erg s-1 for the strongest ten lines of each Au ion species are presented, and in Table 3 the corresponding values for the three Pt ions. Again the photon luminosities are computed with a reference mass of M⊙, an electron temperature = 0.15 eV and an electron density of = 106 cm-3. Clearly the luminosity strengths, for all six ions of both Au and Pt, are orders of magnitude lower than those predicted for W in Table 1, but the strongest transitions are similarly those among the low lying levels. For completeness in Figures 3 and 4 the corresponding broadened luminosity density profiles as a function of wavelength are plotted, to visually display these strong features. Given the lack of discernible features at these wavelengths at the late times of AT2017gfo and AT2023vfi for Pt and Au, this places weak constraints on the masses of these elements as previously stated by (Gillanders et al., 2021). It is therefore the case that higher quality observed data is required to make more conclusive estimates of the presence of the elements in future KN events, in addition to accurate theoretical data for representative species.
@@ 3.2 Ion-mass quantification for W
The late-time-Spitzer (Villar et al., 2018; Kasliwal et al., 2019) observation of AT2017gfo and the observation of AT2023vfi (Levan et al., 2024) have produced interest in a potential feature at around m. The ions W iii and Se iii, which both have thermally accessible fine structure lines at around this wavelength. A first analysis was performed by Hotokezaka et al. (2022) using distorted-wave collision data calculated using the hullac code (Bar-Shalom et al., 2001). Here, a revised analysis of the parameter space for W iii using the newly published -matrix collisional data of McCann et al. (2024) for AT2017gfo is presented, and additionally a new analysis for AT2023vfi. This new atomic data is expected to be more accurate at low temperatures due to the inclusion of resonances when calculating the collision strengths.
@@
For AT2023vfi, the feature has been integrated to have a specific luminosity of erg s-1 (Gillanders & Smartt, 2024). Enforcing this luminosity, and a late time electron density of cm-3 (Levan et al., 2024), at an electron temperature of 3000 K (Levan et al., 2024), the -matrix data of McCann et al. (2024) requires a W iii mass of M⊙. As for AT2017gfo, this calculation is based on the sum of the 4.43 and 4.54 m transitions. Given a total ejecta mass of M⊙ (Levan et al., 2024), this mass of W iii represents a fraction of around %. Without a quantitative treatment of ionization fraction, an elemental mass is difficult to quantify. Assuming reasonably similar fractions for the near neutral stages of W, it is perhaps feasible that up to % could potentially be present in the ejecta.
Figure 5 shows the synthetic W spectrum compared with the JWST spectrum of AT2023vfi (Levan et al., 2024). The centroid and width of the m feature was fitted by Gillanders & Smartt (2024), which is also shown in Figure 5. The spectrum was produced using an electron temperature of 0.26 eV and an electron density of cm-3. A Gaussian kernel was again assumed with a full-width-half-maximum of (Gillanders & Smartt, 2024). With a W iii mass of M⊙, a good match to the JWST data can be found assuming a continuum of a power law added to a blackbody at 620K. This blackbody temperature is in rough agreement with those suggested by Gillanders & Smartt (2024) and Levan et al. (2024).
@@
With increasing temperature at fixed density it is naturally found that less mass of W iii is required to produce the feature. The green dashed line on the top panel of Figure 6 shows the implications of a temperature of 0.86 eV (temperatures this high have been suggested by the models of Pognan et al., 2022). In this case, a relatively low mass of M⊙ of W iii is required to produce the feature. However, at such temperatures significant emission should begin to appear in the m range from W ii and iii and at around m from W i, as shown in Figure 7. While this in principle could constrain the mass of W i, the excess emission expected from W iii that is not present in the observation perhaps prevents this from being a valid identification. In summary, for W iii to be responsible for the excess flux at around m at the rough density regime of cm-3, it is required that the electron temperature also remain relatively low at around say 2500 K. Conversely, at low electron temperatures - say at around the K blackbody temperatures reported by Gillanders & Smartt (2024) and Levan et al. (2024), the feature requires unreasonably high masses of around M⊙ of W iii.
@@ 4 Implications of the limits on the ejected mass of W
In this section, we compare theoretical merger models based on hydrodynamical simulations and nuclear network calculations with the estimated masses of W iii derived in the previous section, i.e. for AT2017gfo and for AT2023vfi. Observational uncertainties of these estimates are discussed in the previous section, e.g. assumptions about the temperature and the possibility of line blending. The merger models only provide the total amount of W without distinguishing different ionization states and are themselves connected with uncertainties, some of which are briefly addressed below. Despite these uncertainties this comparison reveals a broad consistency between theoretical predictions and the observations and highlights the particular usefulness of W measurements to constrain the properties of the merger outflow and the abundance of other elements.
@@
In Table 4 we list the results from a selection of neutron-star merger models, drawn from Just et al. (2023) and Sneppen et al. (2024), which strive for a complete and consistent description of all mass ejection channels. The set consists of four models taken from Just et al. (2023), two for equal-mass binaries ("sym") and two with a binary mass ratio of ("asy"). In order to capture the effects of turbulent angular-momentum transport during the post-merger evolution, these models adopt a viscosity model (Shakura & Sunyaev, 1973). For each of the aforementioned two cases we consider one model with a fairly strong (“n1-a6”) and one model with relatively weak (“n10-a3”) viscosity. The models of Just et al. (2023) only involve relatively long-lived neutron-star remnants that collapse 100 ms after merger or later (cf. in Table 4). We supplement these with two models (one equal mass, one asymmetric) from Sneppen et al. (2024) having short-lived remnants (i.e. collapsing at 10 ms after merger, labelled "short"; see references for details). The simulations consider neutron-stars with masses comparable to those in AT2017gfo. The total amount of ejecta in these simulations is also broadly consistent with the estimated total ejecta mass of AT2017gfo, namely about 0.03–0.05 (e.g. Smartt et al., 2017). In addition to the initial binary parameters and merger dynamics, nucleosynthesis yields are also affected by nuclear-physics uncertainties, which are particularly significant for heavy, neutron-rich nuclei. To explore this sensitivity, we add two further models that adopt the DZ31 nuclear mass model (Duflo & Zuker, 1995) compared to the models in Just et al. (2023); Sneppen et al. (2024), which adopted the HFB21 model (Goriely et al., 2010) as well as the consistently derived and rates in Mendoza-Temis et al. (2015). These two models are chosen to have all other simulation parameters matching previous cases such that the influence of changing the nuclear physics input alone can be ascertained.
@@
Elemental abundances in a KN are generally time dependent, because the radioactive decay half-lives are comparable to the relevant observable timescales of hours to months. The dominating decay mode, beta-decay, maintains constant the isotopic abundances but modifies the elemental abundances. To exemplify this, we show in Figure 8 the time evolution of the relative W mass fraction resulting in the simulation “sym-n1-a6” with the HFB21 nuclear mass model along with the contributions from individual isotopes. At around 1 day, 187W decays, and the total mass fraction decreases. The decays of the other two isotopes 188W and 185W (with half lives of around 70 and 75 days, respectively) lead to a further reduction (see supplemental material of Wu et al., 2019), despite a minor compensation from the decays of 184Ta and 183Ta. As a consequence, the mass fraction of W decreases by a factor of from the first hours to around two months after merger. Notice that this decrease should occur in any r-process model that reproduces the isotopic abundance pattern observed in the solar system around the third r-process peak. Given that the observations of AT2017gfo and AT2023vfi discussed in Section 3.2 were taken at 43 days and 29 days, respectively, after the merger, we base the following quantitative comparison on the mass fractions taken at 1 month, which are listed in Table 4.
@@
The predicted W masses scatter in the different merger models between M⊙ and M⊙ and are thus broadly consistent with the mass of W iii derived through the luminosity of the 4.5 feature in the spectra of AT2017gfo and AT2023vfi , providing tentative evidence that the inference presented in Section 3.2 yields reasonable estimates. The agreement is particularly good for AT2017gfo. Although most of the models give W masses that are a factor of 2–3 smaller than our estimated M⊙ (cf. Section 3.2), the full spread within the models encompasses this value. Also, several of the simulations yield mass fractions for W, (W), that are close to the estimated fraction of 0.3 % (Section 3.2). The short-lived merger models tend to yield somewhat higher values of (W) in slightly better agreement with the mass-fraction estimate from Section 3.2, however, the total ejecta masses in these models are slightly too small to be compatible with AT2017gfo. Considering AT2023vfi, the inferred W iii mass of M⊙ is up to one order of magnitude in excess of what the simulations predict, which were, however, originally set up to reproduce AT2017gfo-like systems. At this point we cannot judge if this discrepancy is a deficiency of the approach adopted in this study keeping in mind, in particular, as noted in Section 3.2, that the ion mass inferred from observations may be an overestimate if line blending is occurring, or only an upper limit if the feature is dominated by another process. Alternatively, this discrepancy may originate from different physical properties of the outflow in AT2023vfi compared to AT2017gfo. As discussed below, a higher W mass points to more neutron-rich ejecta, which may result from a binary with different masses than AT2017gfo. Since the gravitational-wave signal of AT2023vfi was not detected, the binary masses in this event are unknown. High ejecta masses and very neutron-rich outflows may also result from a neutron-star black-hole merger (e.g. Kyutoku et al., 2021), which may also be a possible progenitor of AT2023vfi. The numbers in Table 4 also show very clearly that the nuclear physics input for the network calculation has a significant impact on the W production, with DZ31 yielding systematically higher W masses (compare columns 2 and 8 or columns 7 and 9, respectively). Other aspects such as beta decay and fission rates may also impact the final abundance yields. The uncertainties of the nuclear model may thus represent another possible contribution explaining the difference between the theoretically predicted abundance and the W iii mass inferred from observations. Obviously, the current set of merger simulations is not exhaustive, and other physical ingredients (e.g. a different nuclear equation of state in the simulation) may yield different W masses. Moreover, due to their high level of complexity, the hydrodynamical simulations still adopt a number of approximations (e.g. in the neutrino transport and treatment of small-scale turbulence) and may not be fully resolved numerically.
@@
The robust pattern for the ratio of the W mass to the total ejecta mass with below 0.2, , suggests the possibility to use a measured W mass as a proxy for the amount of neutron-rich ejecta, the theoretical prediction of which, however, is still affected by nuclear-physics uncertainties. Based on the values from Table 4, neutron-rich material should roughly amount to about 40 times (14 times) the inferred W mass for HFB21 (DZ31), i.e. to M⊙ (M⊙) for AT2017gfo and to M⊙ (M⊙) for AT2023vfi. For this neutron-rich material, the measured W mass constraints the production of non thermal electrons by beta decay. We find that for times later than a month 188W represents the dominating contribution.
@@
As an estimate of the uncertainties in the analysis and interpretation, which are typically a factor of a few (see Section 3.2), in Figure 10 we show a green band corresponding to a factor of two variation in both directions and a horizontal green band illustrating the corresponding uncertainties in (using the same conversion as above). Despite the remaining uncertainties our study suggests that a significant amount of lanthanides was produced in AT2017gfo and AT2023vfi. We note that our estimated range of the lanthanide mass fraction (neglecting the small contributions from actinides) is in good agreement with the mass fraction of lanthanides that, according to the analysis of Ji et al. (2019), would be required by the observations of metal-poor stars for mergers to represent the main source of -process elements. Interestingly, the high lanthanide mass fractions that we find exceed most of the estimates made based on the light curve of AT2017gfo in the literature surveyed by Ji et al. (2019). Similarly, our work predicts that a sizable amount of third-peak -process elements of about was co-produced (see lower panel of Figure 10).